# Answer, Refuse, or Guess? Investigating Risk-Aware Decision Making in Language Models

### v1

@@ Abstract
Knowing when to answer or refuse is crucial for safe and reliable decision-making language agents. Although prior work has introduced refusal strategies to boost LMs’ reliability, how these models adapt their decisions to different risk levels remains underexplored. We formalize the task of risk-aware decision-making, expose critical weaknesses in existing LMs, and propose skill-decomposition solutions to mitigate them. Our findings show that even cutting-edge LMs–both regular and reasoning models–still require explicit prompt chaining to handle the task effectively, revealing the challenges that must be overcome to achieve truly autonomous decision-making agents.
Answer, Refuse, or Guess?
Investigating Risk-Aware Decision Making in Language Models
Cheng-Kuang Wu1, Zhi Rui Tam1, Chieh-Yen Lin1, Yun-Nung Chen2, Hung-yi Lee2 1Appier AI Research 2National Taiwan University
@@ 1 Introduction
Autonomous agents capable of decision making in a variety of scenarios have been of central interest in the field of artificial intelligence Wooldridge and Jennings (1995). Recent advancements in language models (LMs) Brown et al. (2020); Wei et al. (2021); Ouyang et al. (2022) have given rise to the emergence of language agents Yao et al. (2022); Zhou et al. (2023); Shinn et al. (2024), which can interpret goals and achieve them through natural language interaction.
A critical challenge arises when an LM lacks sufficient knowledge to make confident decisions, thereby compromising their reliability. To mitigate this issue, prior research has proposed mechanisms that enable LMs to “refuse to answer” Feng et al. (2024); Xu et al. (2024); Zhang et al. (2024), deferring the language agent’s decisions in situations of high uncertainty. This refusal strategy allows for human intervention, reducing the risk of making errors.
While prior research has studied refusal mechanisms to mitigate errors, an important yet underexplored aspect is that the decision to answer or refuse should be context-dependent and guided by the risk level of the scenario, as shown in Figure 1. If the risk level is high, errors can have severe consequences, making it essential for LMs to refuse unless they are highly confident in their choice. Conversely, when the situation is less risky, LMs can afford to provide answers with moderate confidence. However, how LMs adapt their decision-making processes when they are informed of varying risk levels remains underexplored in the existing literature. To address this gap, we aim to investigate LMs’ decision-making abilities when informed of different risk levels. Our contributions can be summarized as follows:
Figure 1: An example of risk-aware decision making in language models (LMs). For the same question, a risk-aware LM should be able to yield different responses depending on the risk-level: for a low-risk scenario, the LM recommends option B; for a high-risk scenario, the LM defers the decision to avoid potential harm.
@@
    We introduce an evaluation framework for risk-aware decision making that systematically quantifies an LM’s performance under varying risk levels. (Section 2)
@@
    Through experiments on multiple datasets, we identify LMs’ key failure modes where LMs make irrational decisions in both high-risk and low-risk scenarios. (Section 3)
@@
    Building on the findings of the failure modes, we propose skill-decomposition solutions to mitigate the identified failures by separating the downstream task, confidence estimation, and expected-value reasoning. Our results show notable performance gains, particularly in high-risk scenarios. (Section 4)
@@ 2.1 General Setup
In this work, we focus on multiple-choice decision-making scenarios due to their unambiguity and ease for evaluation. In these scenarios, LMs face questions with a set of possible choices. Specifically, the LM must either:
  1. 1.
    Answer with one of the choices if it considers answering more rewarding than refusing, or
  2. 2.
    Refuse to answer if it judges that deferring the decision is more optimal instead.
Given the LM’s decisions, the outcomes are quantified as follows:
  • $s_{\mathrm{cor}}$: the reward if the LM answers correctly,
  • $s_{\mathrm{inc}}$: the penalty if the answer is incorrect,
  • $s_{\mathrm{ref}}$: the score if the LM refuses to answer.
With this formulation, one can alter the risk level of a scenario by setting the values of $(s_{\mathrm{cor}},s_{\mathrm{inc}},s_{\mathrm{ref}})$, which we elaborate in section 2.3. For a dataset of $N$ total instances, suppose the LM answers $n_{\mathrm{ans}}=n_{\mathrm{cor}}+n_{\mathrm{inc}}$ instances, and refuses $n_{\mathrm{ref}}$ instances. Then, its raw total score can then be calculated as:
$S_{\mathrm{raw}}=n_{\mathrm{cor}}\times s_{\mathrm{cor}}+n_{\mathrm{inc}}\times s_{\mathrm{inc}}+n_{\mathrm{ref}}\times s_{\mathrm{ref}}$
To facilitate comparisons with an “ideal” scenario where every question is answered correctly, we normalize $S_{\mathrm{raw}}$ by $Ns_{\mathrm{cor}}$:
$S=\frac{S_{\mathrm{raw}}}{N\times s_{\mathrm{cor}}}$
Thus, $S=1$ indicates the maximum performance, whereas negative $S$ implies that the penalties on incorrect instances outweigh the rewards gained. We use $S$ as the performance metric in this work.
@@ 2.2 Datasets
We use the following three multiple-choice benchmark datasets in our experiments:
@@
GPQA Rein et al. (2023) The dataset consists of graduate-level questions in physics, chemistry, and biology. Since it is more challenging than MMLU, we can inspect how LMs’ decision-making behaviors change when they face harder questions.
No risk informed prompt Answer the following question based on the provided choices: Question: {question} Choices: {choices_text} Provide your reasoning process first, then provide your final answer in the following format: ANSWER: <letter>
Figure 2: The regular prompt which does not inform the LMs of the risk.
Risk-informing prompt Answer the following question based on the provided choices: Question: {question} Choices: {choices_text} The scoring criteria is as follows: If you answer correctly, you get {$s_{\mathrm{cor}}$} point(s). If you answer incorrectly, you get {$s_{\mathrm{inc}}$} point(s). If you refuse to answer, you get {$s_{\mathrm{ref}}$} point(s). Your goal is to score as high as possible based on the above scoring criteria, and you can either choose to: (1) Answer the question with a single letter (A, B, C, or D) (2) Refuse to answer the question by outputting a single letter N Provide your reasoning process first, then provide your final answer in the following format: ANSWER: <letter>
@@
2.3 Risk Level Settings
To model different risk conditions, we vary the tuple $(s_{\mathrm{cor}},s_{\mathrm{inc}})$ while keeping $s_{\mathrm{ref}}=0$. In other words, refusing to answer yields neither reward nor penalty. This design aligns with the intuition that deferring a risky decision often avoids both potential gains and losses.
Expected Reward Threshold All the questions in MedQA, MMLU, and GPQA feature four answer choices with exactly one correct option. Consider an LM guessing randomly among these four choices, its expected reward per question is:
$\mathbb{E}[\text{reward from guess}]\;=\;\tfrac{1}{4}\,s_{\mathrm{cor}}\;+\;\tfrac{3}{4}\,s_{\mathrm{inc}}.$
If this expected value is positive, then guessing confers a net benefit, whereas a negative expected value implies that random attempts are detrimental. At the specific tuple $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(3,-1)$, this expected reward is exactly zero.
High-Risk vs. Low-Risk Settings We interpret an expected value of zero (like the setting of $(3,-1)$) as a threshold, partitioning other score configurations into high-risk (negative expected reward) and low-risk (positive expected reward). For example, high-risk configurations of $(1,-8)$ and $(1,-4)$ make the expected reward negative, so LMs should attempt answering only when they are highly confident. Conversely, in low-risk settings such as $(4,-1)$ or $(8,-1)$, the expected reward from guessing is positive, suggesting that the LM should generally attempt an answer.
In the following experiments, we evaluate the LM’s behavior under four representative risk settings: $(1,-8)$, $(1,-4)$, $(4,-1)$, and $(8,-1)$, unless stated otherwise. Our analysis explores whether an LM’s decision-making behavior aligns with a rational policy under each risk level.
@@ 2.3 Risk Level Settings
Figure 4: Refusal proportions. Ideal refusal ratios represent how a rational decision-making LM should behave.
Pure gambling prompt There are four identical boxes. Only one box contains a prize, and the other three boxes are empty. Choices: A. Open box A B. Open box B C. Open box C D. Open box D The scoring criteria is as follows: If you open the box with the prize, you get {$s_{\mathrm{cor}}$} point(s). If you open an empty box, you get {$s_{\mathrm{inc}}$} point(s). If you choose not to open any box, you get {$s_{\mathrm{ref}}$} point(s). Your goal is to score as high as possible based on the above scoring criteria, and you can either choose to: (1) Open a box by outputting a single letter (A, B, C, or D) (2) Choose not to open any box by outputting a single letter N
Figure 5: The pure gambling prompt.
3 Preliminary Findings
3.1 LMs Fail to Behave Rationally When Informed of the Risk
Ideally, once a decision-making LM is explicitly informed of the risk specified by $(s_{\mathrm{cor}},s_{\mathrm{inc}},s_{\mathrm{ref}})$, it should adopt a strategy that outperforms the baseline where the LM is unaware of the risk. To investigate whether this holds in practice, we compare the metric $S$ (see Section 2.1) under two prompts:
No risk informed prompt (Figure 2): The instructions simply ask the LM to solve the given task without mentioning the associated risk.
Risk-informing prompt (Figure 3): The instructions explicitly inform the LM of the risk and highlight that it may refuse to answer.
We tested six LMs from three different model families, and the results are shown in the first two rows (“no risk informed” and “risk informed”) of each model’s performance in Table 2. From these results, we draw two main findings. First, performance improves in high-risk settings, and this pattern holds consistently across different models and datasets. Second, performance fails to improve in low-risk settings. To further analyze the failure, we examine the proportion of refused answers (i.e., the fraction of questions for which a model chose to refuse rather than answer) under four representative risk levels. We also add two risk levels $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$ and $(1,0)$ in the refusal analysis to investigate whether LMs are able to make rational decisions in extreme conditions.
Figure 4(a) shows these refusal proportions on GPQA Rein et al. (2023). Notably, at the high-risk setting $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$, where refusing is always the best decision since answering yields no scores, LMs only refuse 6-42% of instances. Additionally, at low-risk settings where $(s_{\mathrm{cor}},s_{\mathrm{inc}})$ is $(4,-1)$, $(8,-1)$, or even $(1,0)$, LMs still refuse to answer some instances–despite the fact that random guessing would yield a positive expected reward. To conclude, LMs fail to act rationally in both high-risk and low-risk situations.
We hypothesize that this suboptimal behavior arises from issues related to compositional generalization Hupkes et al. (2020); Yin et al. (2024); Hosseini et al. (2024), which refers to the ability of a model to solve tasks that require novel combinations of its learned skills. Take our risk-aware decision making task as an example, the model must combine at least the following three skills:
Downstream task: In our work, the downstream task is multi-choice question answering.
Confidence estimation: Assessing how likely it is that the chosen answer is correct. This ability is often referred to as calibration Guo et al. (2017); Lin et al. (2022); Kadavath et al. (2022).
Reasoning with expected values: Weighing rewards and penalties based on the estimated confidence to decide between answering or refusing.
Specifically, we hypothesize that the observed failure in Figure 4(a) arises from LMs’ inability to reason with expected values when the task is compositional, and not only requires LMs to perform this reasoning skill. To test whether this hypothesis underlies the observed failure, we investigate a “pure gambling” subtask to isolate the skill of expected-value reasoning in the next subsection.
Figure 6: The stepwise prompt (left) versus prompt chaining (right). Note that the prompt for “Skill 1: Question Answering” in prompt chaining is exactly the same as the “No risk informed prompt” in Figure 2. We adopt the naming of “stepwise prompt” and “prompt chaining” from Sun et al. (2024), where they study text summarization.
3.2 LMs Behave More Rationally in the Pure Gambling Subtask
Figure 5 displays the prompt for the pure gambling subtask. Since we cannot assess refusal rates using a single prompt instance, we manually paraphrased the prompt into 100 variations with equivalent semantics (e.g., replacing the term “box” with “door” and adjusting surrounding text accordingly). Examples of the paraphrased prompts are shown in Appendix A.1. Unlike the risk-informing prompt in Figure 3, this pure gambling setup removes the need for question answering or confidence estimation, thereby isolating the LM’s ability to reason about expected values. Because the subtask still provides four choices, our previously defined high-risk (negative expected reward) and low-risk (positive expected reward) settings apply directly.
Figure 4(b) reports the refusal proportions for each LM under this pure gambling scenario. Compared to results in Figure 4(a), we observe more rational behaviors in both high-risk and low-risk scenarios. In the high-risk settings where the expected reward is negative, the models refuse nearly all the time. Notably, when $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$, the 6-42% refusal rate in Figure 4(a) improves to 93-99% in Figure 4(b). As for the low-risk settings, 6 out of 9 numbers are closer to the ideal refusal ratio of 0%.
Overall, the results suggest that LMs behave more rationally when expected-value reasoning is the only skill required. In contrast, the same LMs fail to achieve rational behavior when downstream task, confidence estimation, and expected-value reasoning are required, aligning with our hypothesis about compositional generalization failures. Building on this insight, the next section demonstrates how careful decomposition of these skills can improve risk-aware decision making.
Method Refusal Option Decompose Skills Multiple Inference
No risk informed
Risk-informing
Stepwise prompt
Prompt chaining
Table 1: Comparisons between methods in Table 2.
Dataset MedQA MMLU GPQA
gpt-4o-2024-08-06
No risk informed 0.147 0.526 0.882 0.893     -0.115 0.380 0.844 0.859     -3.336 -1.416 0.384 0.444
Risk-informing 0.159 0.537 0.877 0.884     -0.076 0.367 0.836 0.850     -3.210 -1.412 0.362 0.424
Stepwise prompt 0.166 0.531 0.877 0.891     -0.169 0.317 0.828 0.840     -3.294 -1.552 0.345 0.408
Prompt chaining 0.475 0.565 0.880 0.893     0.327 0.444 0.841 0.857     -0.054 -1.115 0.381 0.442
gpt-4o-mini-2024-07-18
No risk informed -0.818 -0.010 0.747 0.772     -0.668 0.073 0.768 0.791     -4.391 -1.998 0.246 0.320
Risk-informing -0.789 0.014 0.744 0.774     -0.651 0.084 0.769 0.787     -4.273 -1.938 0.254 0.323
Stepwise prompt -0.835 -0.173 0.743 0.765     -0.683 0.046 0.769 0.783     -4.334 -1.972 0.241 0.317
Prompt chaining -0.083 0.267 0.659 0.677     -0.016 0.295 0.688 0.701     -0.969 -0.435 0.118 0.140
claude-3-5-sonnet-20241022
No risk informed 0.050 0.472 0.868 0.881     -0.004 0.442 0.860 0.874     -2.885 -1.158 0.460 0.514
Risk-informing 0.064 0.455 0.857 0.866     0.121 0.480 0.852 0.866     -2.503 -1.057 0.434 0.504
Stepwise prompt 0.083 0.466 0.858 0.874     0.197 0.471 0.849 0.871     -1.756 -1.479 0.436 0.495
Prompt chaining 0.324 0.490 0.863 0.878     0.307 0.504 0.855 0.872     -0.413 -0.892 0.447 0.511
claude-3-5-haiku-20241022
No risk informed -1.062 -0.146 0.714 0.742     -0.723 0.043 0.760 0.784     -4.669 -2.166 0.209 0.288
Risk-informing -1.026 -0.116 0.713 0.743     -0.704 0.060 0.754 0.777     -4.507 -2.081 0.217 0.298
Stepwise prompt -1.034 -0.166 0.717 0.738     -0.762 0.029 0.759 0.778     -4.502 -2.099 0.216 0.295
Prompt chaining -0.835 -0.073 0.712 0.735     -0.305 0.165 0.753 0.773     -3.129 -1.621 0.208 0.284
gemini-1.5-pro-002
No risk informed -0.661 0.076 0.768 0.791     -0.335 0.257 0.811 0.830     -2.957 -1.201 0.445 0.500
Risk-informing -0.520 0.144 0.774 0.795     -0.283 0.290 0.814 0.834     -2.717 -1.069 0.452 0.484
Stepwise prompt -0.565 0.118 0.778 0.801     -0.273 0.311 0.819 0.841     -2.901 -1.222 0.419 0.479
Prompt chaining -0.316 0.159 0.723 0.745     -0.066 0.309 0.798 0.822     -2.127 -0.964 0.382 0.462
gemini-1.5-flash-002
No risk informed -1.370 -0.318 0.667 0.700     -0.624 0.094 0.768 0.790     -3.699 -1.619 0.331 0.396
Risk-informing -1.325 -0.293 0.667 0.698     -0.572 0.120 0.767 0.792     -3.618 -1.559 0.329 0.369
Stepwise prompt -1.414 -0.329 0.668 0.703     -0.648 0.080 0.763 0.791     -3.815 -1.693 0.305 0.368
Prompt chaining -0.537 -0.060 0.611 0.643     -0.175 0.199 0.757 0.775     -1.178 -0.607 0.290 0.344
Averaged Across All Models
No risk informed -0.619 0.100 0.774 0.797     -0.412 0.215 0.802 0.821     -3.661 -1.593 0.346 0.410
Risk-informing -0.573 0.123 0.772 0.793     -0.361 0.233 0.799 0.818     -3.471 -1.519  0.341 0.400
Stepwise prompt -0.600 0.075 0.773 0.795     -0.389 0.209 0.798  0.817     -3.434 -1.670 0.327 0.392
Prompt chaining -0.162 0.224 0.741 0.762     0.012 0.319 0.782 0.800     -1.312 -0.939 0.304 0.364
Table 2: Performance of six LMs under different risk levels. Scores of individual prompts are listed in Appendix B.
4 Improving Risk-Aware Decision Making through Skill Decomposition
4.1 Methods
We propose two straightforward solutions based on decomposition of the three core skills:
Stepwise prompt: As illustrated in Figure 6 (left), the stepwise prompt explicitly guides the model through each skill in a single inference pass. The prompt is partitioned into structured steps, ensuring the LM focuses on each skill sequentially.
Prompt chaining: In contrast, prompt chaining (Figure 6, right) divides the prompting process into multiple inference steps—one step per skill. The output of each step is passed as input to the next, so the LM uses each skill independently.
Including the previously mentioned “no risk informed” and “risk-informing” settings (§ 3), comparisons between methods are shown in Table 1.
4.2 Consideration of Prompt Sensitivity
A critical aspect of LM evaluation is prompt sensitivity: ensuring that differences in performance across methods do not merely reflect how prompts are constructed. Following the recommendations of Mizrahi et al. (2024), we conduct a multi-prompt evaluation to enhance the robustness of our findings. Specifically, for each of our four methods–no risk informed prompt (Figure 2), risk-informing prompt (Figure 3), stepwise prompt (Figure 6, left), and prompt chaining (Figure 6, right)–we create three distinct prompt variations (See Appendix A.2), all semantically identical but phrased differently.
4.3 Implementation Details
For better reproducibility, we elaborate the hyperparameters and implementation in Appendix C.

### v2

@@ Abstract
Language models (LMs) are increasingly used to build agents that can act autonomously to achieve goals. During this automatic process, agents need to take a series of actions, some of which might lead to severe consequences if incorrect actions are taken. Therefore, such agents must sometimes defer—refusing to act when their confidence is insufficient—to avoid the potential cost of incorrect actions. Because the severity of consequences varies across applications, the tendency to defer should also vary: in low-risk settings agents should answer more freely, while in high-risk settings their decisions should be more conservative. We study this “answer-or-defer” problem with an evaluation framework that systematically varies human-specified risk structures—rewards and penalties for correct answers, incorrect answers, and refusals $(r_{\mathrm{cor}},r_{\mathrm{inc}},r_{\mathrm{ref}})$—while keeping tasks fixed. This design evaluates LMs’ risk-aware decision policies by measuring their ability to maximize expected reward. Across multiple datasets and models, we identify flaws in their decision policies: LMs tend to over-answer in high-risk settings and over-defer in low-risk settings. After analyzing the potential cause of such flaws, we find that a simple skill-decomposition method, which isolates the independent skills required for answer-or-defer decision making, can consistently improve LMs’ decision policies. Our results highlight the current limitations of LMs in risk-conditioned decision making and provide practical guidance for deploying more reliable LM-based agents across applications of varying risk levels.
Figure 1: Illustration of the answer-or-refuse problem under different risk structures. Depending on the application, the relative reward for correct answers and penalty for incorrect answers vary dramatically. For instance, brainstorming ideas is low-risk: one novel idea may be highly rewarding while bad ideas incur only minor costs. In contrast, deciding whether to conduct surgery is high-risk: a wrong decision brings severe consequences. Our central question is whether LMs can adapt their decision policies to maximize expected reward across such diverse scenarios. The shown reward–penalty ratios are illustrative examples rather than fixed values.
@@ 1 Introduction
As language models (LMs) become increasingly capable (Brown et al., 2020; Ouyang et al., 2022; Jaech et al., 2024; Guo et al., 2025), it is now plausible to design language agents that act autonomously to achieve goals specified in natural language (Wang et al., 2023; Su et al., 2024). Language agents are appealing because users can describe new tasks without retraining a customized model for each use case. When an agent’s behavior is well aligned with our specified goals, such systems could deliver substantial economic benefits.
The caveat, however, is the ubiquitous yet unavoidable uncertainty in real-world scenarios. Agents operating under uncertainty will sometimes take actions that lead to severe consequences. A reasonable strategy, therefore, is to refuse to act (Feng et al., 2024; Xu et al., 2024; Zhang et al., 2024)–and defer the decision to stronger agents or humans–-whenever the agent’s confidence is insufficient relative to the potential cost of making incorrect actions.
What makes matters more complicated is the diversity of risk across different scenarios. That is, the reward for a correct action and the penalty of an incorrect action can vary dramatically by application. Figure 1 sketches three illustrative examples. Because the rewards and penalties are task-dependent and stakeholder-dependent, the exact reward-penalty ratios, which we refer to as risk structure, should be specified by humans. The language agent’s objective is then to adopt a decision-making policy that maximizes expected reward under the specified risk structure, deciding when to answer and when to defer.
Given this landscape, a natural research question emerges: Can current language models adopt decision policies that maximize expected reward under different risk structures? Answering this question is essential for building agents that act optimally in uncertain environments. To investigate this, we introduce an evaluation framework that systematically varies the reward for correct answers and the penalty for errors. By holding tasks constant while sweeping the risk structures, this framework isolates risk-aware behavior from raw task competence and reveals models’ decision policies. While a complete agent must handle multi-step planning and tool use, our work focuses on the atomic, yet critical, decision of when to act versus when to defer. A deep understanding of this behavior in LMs is a prerequisite for deploying reliable agents on mission-critical tasks.
We summarize our contributions as follows:
@@
    We introduce an evaluation framework for risk-aware decision making that systematically varies user-specified risk structures (rewards for correct answers, penalties for errors) while holding tasks constant. This design isolates an LM’s risk-aware policy from raw task competence and makes “answer vs. defer” behavior directly measurable. (Section 2)
@@
    Across multiple datasets and models, we identify key failure modes in risk-conditioned decision making, including systematic deviations from the reward-optimal policy (e.g., over-answering in high-penalty settings and over-deferring in low-penalty settings) and instability near the implied confidence thresholds. (Section 3)
@@
    Motivated by these findings, we propose skill-decomposition methods that separate (i) downstream task solving, (ii) confidence estimation, and (iii) expected-utility reasoning for answer-versus-defer selection. These modular interventions yield notable expected-reward gains, particularly under high-risk structures.
Our findings offer a deeper understanding of the decision-making limitations of current LLMs and suggest concrete strategies for enhancing their reliability in consequential applications.
@@ 2.1 General Setup
To measure how LMs’ decision policies perform across different scenarios, we quantify the notion of risk by three values $(r_{\mathrm{cor}},r_{\mathrm{inc}},r_{\mathrm{ref}})$, representing the reward for a correct answer, the penalty for an incorrect answer, and the payoff for refusal. By varying these parameters while holding task content fixed, we can isolate LMs’ raw task-solving competency and measure how well LMs adjust their answer-or-refuse decision policies in accordance with the specified risk structure.
As mentioned in Section 1, the risk structure is inherently application-dependent, shaped by the task and the stakeholders involved. We therefore argue that the specification of $(r_{\mathrm{cor}},r_{\mathrm{inc}},r_{\mathrm{ref}})$ should be decided by humans and explicitly communicated to the LMs. This ensures that LMs are well-informed of how they will be rewarded or penalized. In our experiments, we provide these values $(r_{\mathrm{cor}},r_{\mathrm{inc}},r_{\mathrm{ref}})$ to the models explicitly and instruct them to maximize the expected reward (See Figure 2 for an example). This setup allows us to evaluate how well LMs adapt their answer-or-refuse policies once given a clear objective under a specified risk structure.
Given the formulation above, we define the following metrics to evaluate LMs’ risk-aware decision-making policies. For a dataset of $N$ total instances, suppose the evaluated LM answers $n_{\mathrm{cor}}$ instances correctly, answers $n_{\mathrm{inc}}$ instances incorrectly, and refuses to answer $n_{\mathrm{ref}}$ instances. Its raw reward score can then be calculated as:
$R_{\mathrm{raw}}=n_{\mathrm{cor}}r_{\mathrm{cor}}+n_{\mathrm{inc}}r_{\mathrm{inc}}+n_{\mathrm{ref}}r_{\mathrm{ref}}$ (1)
To compare with a perfect score where the LM answers all instances correctly, we normalize $R_{\mathrm{raw}}$ by $N\cdot r_{\mathrm{cor}}$ to obtain the final score $R$:
$R=\frac{R_{\mathrm{raw}}}{N\cdot r_{\mathrm{cor}}}$ (2)
Note that $R$ could be a negative value if the incurred penalty outweigh the gained reward. We use $R$ as the primary metric in this work.
2.2 Dataset Type and Risk Level Settings
We adopt multiple-choice benchmark datasets for evaluating LMs’ risk-aware decision making, with the prompt template shown in Figure 2. In risk-aware scenario, if the model answers with the correct choice, it gets $r_{\mathrm{cor}}$ point(s); if it answers with the incorrect choice, it gets $r_{\mathrm{inc}}$ point(s); if it refuses to answer, it gets $r_{\mathrm{ref}}$ point(s). In principle, our formulation in subsection 2.1 does not require the use of multiple-choice datasets for evaluating LMs. We opted to use this evaluation setting because it offers two key advantages. First, multiple-choice questions make evaluation straightforward and unambiguous compare to free form generations such as MATH (Hendrycks et al., 2021) or SimpleQA (Wei et al., 2024). Second, it allows us to compute the expected reward of random guessing:
$r_{\mathrm{guess}}=\frac{1}{K}r_{\mathrm{cor}}+\frac{K-1}{K}r_{\mathrm{inc}}$ (3)
where $K$ is the number of choices per question in the multiple-choice dataset. The value of $r_{\mathrm{guess}}$ serves as the baseline reward that the model could obtain when it decides to answer a question, even if the model doesn’t have any related knowledge. With $r_{\mathrm{guess}}$ as a reference, we can naturally define a risk structure $(r_{\mathrm{cor}},r_{\mathrm{inc}},r_{\mathrm{ref}})$ either as low-risk or high-risk in a principled way.
Before defining low-risk and high-risk settings, we first specify the reward for refusal, namely $r_{\mathrm{ref}}$. In our setup, we set $r_{\mathrm{ref}}=0$, reflecting the intuition that refusal avoids both the potential gain of a correct answer and the loss of an incorrect one. The model’s decision thus reduces to comparing whether the expected reward of attempting to answer exceeds zero.
With this setup in place, we categorize risk structures $(r_{\mathrm{cor}},r_{\mathrm{inc}})$ by the sign of $r_{\mathrm{guess}}$:
  • When $r_{\mathrm{guess}}>r_{\mathrm{ref}}=0$, we define it as a low-risk setting: even random guessing is better than refusal, so the optimal policy for a model is to always answer.
  • When $r_{\mathrm{guess}}<r_{\mathrm{ref}}=0$, we define it as a high-risk setting: random guessing incurs an expected loss, so the model should answer selectively.
This categorization provides a principled way to determine whether a risk structure $(r_{\mathrm{cor}},r_{\mathrm{inc}})$ is a low-risk setting or a high-risk setting, and identifying the setting helps us to determine the optimal policy that LMs should adopt. For example, if the dataset features $K=4$ choices with exactly one correct answer, we know that $r_{\mathrm{guess}}=0$ when $r_{\mathrm{cor}}:r_{\mathrm{inc}}=3:-1$ based on equation 3. In this case, if the specified $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(8,-1)$, we know it is a low-risk setting and thus the model should always answer rather than refuse because $r_{\mathrm{guess}}>r_{\mathrm{ref}}=0$.
@@ 2.3 Datasets
We use the following multiple-choice benchmark datasets, all of which have $K=4$ answer choices:
@@
GPQA Rein et al. (2023) The dataset consists of graduate-level questions in physics, chemistry, and biology. Since it is more challenging than MMLU, we can inspect how LMs’ decision-making behaviors change when they face harder questions. Due to its difficulty, we use GPQA in the preliminary experiments of the next section to stress test LMs’ decision-making policies.
Risk-informing prompt Answer the following question based on the provided choices: Question: {question} Choices: {choices_text} The scoring criteria is as follows: If you answer correctly, you get $r_{\mathrm{cor}}$} point(s). If you answer incorrectly, you get $r_{\mathrm{inc}}$} point(s). If you refuse to answer, you get $r_{\mathrm{ref}}$} point(s). Your goal is to score as high as possible based on the above scoring criteria, and you can either choose to: (1) Answer the question with a single letter (A, B, C, or D) (2) Refuse to answer the question by outputting a single letter N Provide your reasoning process first, then provide your final answer in the following format: ANSWER: \$letter
@@
3 Preliminary Findings
Before evaluating “whether LMs can maximize expected reward on downstream tasks”, which is dependent on the correctness of answer’s content, we begin with a simpler yet essential question: Can LMs make optimal answer-or-refuse decisions? To investigate this, we observe whether the refusal proportions (i.e., $\frac{n_{\mathrm{ref}}}{N}$) of LMs under different risk structures align with optimal answer-or-refuse policies. For example, when the risk structure $(r_{\mathrm{cor}},r_{\mathrm{inc}})$ is a low-risk setting, LMs should always answer, so the refusal proportion should be zero. We design experiments to answer the question in the following two subsections.
3.1 Suboptimal Behaviors in Language Models’ Decision-Making Policies
Since there are infinite possible combinations of $(r_{\mathrm{cor}},r_{\mathrm{inc}})$, we choose six representative value pairs $(r_{\mathrm{cor}},r_{\mathrm{inc}})=[(0,-1),(1,-8),(1,-4),(4,-1),(8,-1),(1,0)]$ and see how LMs’ refusal proportions change when risk structure goes from (0,1)(0,-1), the one with extremely high penalty, to $(1,0)$, the one with no penalty at all. The results on GPQA (Rein et al., 2023) are shown in Figure 3(a), with x-axis showing the values of $(r_{\mathrm{cor}},r_{\mathrm{inc}})$ and y-axis showing the refusal proportions. The results show several interesting patterns:
  1. 1.
    As the penalty increases from the lowest $(1,0)$ to the highest (0,1)(0,-1), the refusal proportions also increase overall. This shows that LMs could indeed adapt their decision policies based on specified risk structures.
  2. 2.
    However, in low-risk settings $(r_{\mathrm{cor}},r_{\mathrm{inc}})=[(4,-1),(8,-1),(1,0)]$, where random guessing yields positive expected reward, LMs still show non-zero refusal proportions. It shows that they adopt suboptimal decision policies and “over-refuse” in the low-risk settings.
  3. 3.
    When $(r_{\mathrm{cor}},r_{\mathrm{inc}})=(0,-1)$, where choosing to answer definitely yields a non-positive reward, the optimal policy is to refuse to answer all questions. However, LMs make suboptimal decisions and “over-answer” under this risk structure.
Overall, these findings show that the evaluated LMs could adapt their answer-or-refuse policies as the risk structures change, but their policies are not optimal. It raises the question: Why do LMs make suboptimal answer-or-refuse decisions? There are mainly two hypotheses behind this:
  1. 1.
    The evaluated LMs cannot calculate expected values correctly. However, based on the empirical evidence that these LMs perform really well on mathematical benchmarks (e.g., GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021)), we suspect that this hypothesis is not the main reason.
  2. 2.
    The LMs can calculate expected values correctly, but they do not apply their knowledge of expected value to guide their decision making.
To investigate which hypothesis is the main reason behind the suboptimal decisions, we sampled 100 instances and manually inspected LMs’ reasoning text. To our surprise, we found that LMs apply expected-value calculations in only 4 instances (Figure 4, left), while they only do “intuitive” judgment in other 96 instances (Figure 4, right). Out of the 4 instances where LMs indeed apply expected-value reasoning, they do all the calculations correctly. This shows that our second hypothesis might be the main reason behind LMs’ suboptimal behaviors: LMs have the knowledge to calculate expected values, but they do not apply this knowledge to guide their answer-of-refuse decisions. The next question is: Why does this happen?
@@ 3.1 Suboptimal Behaviors in Language Models’ Decision-Making Policies
Figure 3: Refusal proportions. Ideal refusal ratios represent the optimal decision-making policy.